# Infinitely many pairs of non-isomorphic elliptic curves sharing the same BSD invariants

### v1

@@ 6.2. Pairs of elliptic curves sharing the BSD invariants, Kodaira symbols and minimal discriminants
Combining Proposition 5.4 and Proposition 6.1, for a prime number $D>0$ such that $D\equiv 1\pmod{8}$, $\left(\frac{D}{3}\right)=-1$, $\left(\frac{D}{5}\right)=-1$, $\left(\frac{D}{13}\right)=1$, $\gcd(D,2\cdot 3\cdot 5\cdot 13)=1$, and $D$ is large enough so that $E_{1}^{D}(\mathbb{Q})_{\text{tor}}$ and $E_{2}^{D}(\mathbb{Q})_{\text{tor}}$ do not have any elements of order $>2$, we have $\text{BSD}(E_{1}^{D}/\mathbb{Q})=\text{BSD}(E_{2}^{D}/\mathbb{Q})$ and the Kodaira symbols of $E_{1}$ and $E_{2}$ are the same. We provide a table of pairs of non-isomorphic elliptic curves that share the same BSD invariants.
@@
Real period $\dfrac{0.209\cdots}{\sqrt{|D|}}$ $\dfrac{0.209\cdots}{\sqrt{|D|}}$
@@
$\Sha(E/\mathbb{Q})$dagger $\dagger$
For $\dagger$, note that $\Sha(E/\mathbb{Q})[l^{\infty}]$ remains unknown, but they agree as abelian groups.

### v2

@@ 6.2. Pairs of elliptic curves sharing the BSD invariants, Kodaira symbols and minimal discriminants
Combining Proposition 5.4 and Proposition 6.1, for a prime number $D>0$ such that $D\equiv 1\pmod{8}$, $\left(\frac{D}{5}\right)=-1$, $\left(\frac{D}{13}\right)=1$, $\gcd(D,2\cdot 3\cdot 5\cdot 13)=1$, and $D$ is large enough so that $E_{1}^{D}(\mathbb{Q})_{\text{tor}}$ and $E_{2}^{D}(\mathbb{Q})_{\text{tor}}$ do not have any elements of order >2>2, we have $\text{BSD}(E_{1}^{D}/\mathbb{Q})=\text{BSD}(E_{2}^{D}/\mathbb{Q})$ and the Kodaira symbols of $E_{1}$ and $E_{2}$ are the same. We provide a table of pairs of non-isomorphic elliptic curves that share the same BSD invariants.
@@
Real period $\dfrac{0.209\cdots}{\sqrt{D}}$ $\dfrac{0.209\cdots}{\sqrt{D}}$
@@
$\Sha(E/\mathbb{Q})$ $\dagger$ $\dagger$
For $\dagger$, note that $\Sha(E/\mathbb{Q})[\ell^{\infty}]$ remains unknown, but they agree as abelian groups by Proposition 5.3 (2).